# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $0053\scriptstyle\pm 0.0092 0.0$$033\scriptstyle\pm 0.0058 0.0$$041\scriptstyle\pm 0.0071 0.0$$21\scriptstyle\pm 0.0036 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) $0070\scriptstyle\pm 0.0121 0.0$$133\scriptstyle\pm 0.0231 0.0$$092\scriptstyle\pm 0.0159 0.0$$46\scriptstyle\pm 0.0080 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5000\scriptstyle\pm 0.0500$ $0.4367\scriptstyle\pm 0.0702$ $0.4649\scriptstyle\pm 0.0545$ $0.3040\scriptstyle\pm 0.0457$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0577}$
L3-FT(QA) + KG $0.5267\scriptstyle\pm 0.0250$ $0.5067\scriptstyle\pm 0.0115$ $0.5163\scriptstyle\pm 0.0141$ $0.3480\scriptstyle\pm 0.0128$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.1000$
G-Retriever $0.4846\scriptstyle\pm 0.0537$ $0.4300\scriptstyle\pm 0.0173$ $0.4540\scriptstyle\pm 0.0126$ $0.2937\scriptstyle\pm 0.0106$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
GALAX $\mathbf{0.5433}\scriptstyle\pm\mathbf{0.0252}$ $\mathbf{0.5433}\scriptstyle\pm\mathbf{0.0252}$ $\mathbf{0.5433}\scriptstyle\pm\mathbf{0.0252}$ $\mathbf{0.3733}\scriptstyle\pm\mathbf{0.0238}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
Table 25: Model performances on NB
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0085\scriptstyle\pm 0.0148 0.0$$67\scriptstyle\pm 0.0289 0.011$$13\scriptstyle\pm 0.0196 0.0$$57\scriptstyle\pm 0.0100 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $\mathbf{0.5597}\scriptstyle\pm\mathbf{0.0479}$ $0.5233\scriptstyle\pm 0.0651$ $0.5407\scriptstyle\pm 0.0559$ $0.3718\scriptstyle\pm 0.0523$ $0.8667\scriptstyle\pm 0.2309$ $0.8667\scriptstyle\pm 0.1528$
L3-FT(QA) + KG $0.5507\scriptstyle\pm 0.0704$ $0.5067\scriptstyle\pm 0.0723$ $0.5275\scriptstyle\pm 0.0703$ $0.3603\scriptstyle\pm 0.0667$ $0.9333\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
G-Retriever $0.4322\scriptstyle\pm 0.0444$ $0.3933\scriptstyle\pm 0.0404$ $0.4119\scriptstyle\pm 0.0423$ $0.2599\scriptstyle\pm 0.0341$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.0577}$
GALAX $0.5533\scriptstyle\pm 0.0115$ $\mathbf{0.5533}\scriptstyle\pm\mathbf{0.0115}$ $\mathbf{0.5533}\scriptstyle\pm\mathbf{0.0115}$ $\mathbf{0.3825}\scriptstyle\pm\mathbf{0.0110}$ $0.8667\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
Table 26: Model performances on MESO
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0190\scriptstyle\pm 0.0201 0.0$$167\scriptstyle\pm 0.0208 0.0$$133\scriptstyle\pm 0.0121 0.0$$67\scriptstyle\pm 0.0061 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0192\scriptstyle\pm 0.0228 0.0$$100\scriptstyle\pm 0.0000 0.0$$104\scriptstyle\pm 0.0052 0.0$$53\scriptstyle\pm 0.0026 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5182\scriptstyle\pm 0.0806$ $0.4833\scriptstyle\pm 0.0874$ $0.4995\scriptstyle\pm 0.0808$ $0.3354\scriptstyle\pm 0.0720$ $0.9333\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.1528$
L3-FT(QA) + KG $0.5417\scriptstyle\pm 0.0419$ $0.5233\scriptstyle\pm 0.0416$ $0.5322\scriptstyle\pm 0.0402$ $0.3633\scriptstyle\pm 0.0374$ $0.8667\scriptstyle\pm 0.1155$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
G-Retriever $0.4721\scriptstyle\pm 0.0072$ $0.3733\scriptstyle\pm 0.0462$ $0.4161\scriptstyle\pm 0.0324$ $0.2630\scriptstyle\pm 0.0255$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8000\scriptstyle\pm 0.0000$
@@
Table 27: Model performances on LIHC
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0031\scriptstyle\pm 0.0027 0.0$$067\scriptstyle\pm 0.0058 0.0$$043\scriptstyle\pm 0.0037 0.0$$21\scriptstyle\pm 0.0019 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.4239\scriptstyle\pm 0.0343$ $0.3967\scriptstyle\pm 0.0513$ $0.4096\scriptstyle\pm 0.0435$ $0.2582\scriptstyle\pm 0.0341$ $0.7333\scriptstyle\pm 0.1155$ $0.7000\scriptstyle\pm 0.1732$
L3-FT(QA) + KG $0.4552\scriptstyle\pm 0.0568$ $0.4467\scriptstyle\pm 0.0635$ $0.4509\scriptstyle\pm 0.0600$ $0.2924\scriptstyle\pm 0.0512$ $0.8000\scriptstyle\pm 0.0000$ $0.7667\scriptstyle\pm 0.0577$
G-Retriever $0.3598\scriptstyle\pm 0.0387$ $0.3000\scriptstyle\pm 0.0520$ $0.3270\scriptstyle\pm 0.0467$ $0.1961\scriptstyle\pm 0.0340$ $0.8000\scriptstyle\pm 0.0000$ $0.7667\scriptstyle\pm 0.1155$
@@
Table 28: Model performances on LAML
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $1576\scriptstyle\pm 0.1412 0.0$$200\scriptstyle\pm 0.0173 0.0$$354\scriptstyle\pm 0.0307 0.0$$0182\scriptstyle\pm 0.0158$ $0.1333\scriptstyle\pm 0.2309$ $0.2000\scriptstyle\pm 0.1732$
L3-FT(Med) $1136\scriptstyle\pm 0.0630 0.0$$400\scriptstyle\pm 0.0173 0.0$$495\scriptstyle\pm 0.0008 0.0$$0254\scriptstyle\pm 0.0004$ $0.2000\scriptstyle\pm 0.3464$ $0.1333\scriptstyle\pm 0.1528$
L3-FT(Med) + KG $1085\scriptstyle\pm 0.0867 0.0$$267\scriptstyle\pm 0.0058 0.0$$377\scriptstyle\pm 0.0221 0.0$$0193\scriptstyle\pm 0.0114$ $0.2000\scriptstyle\pm 0.3464$ $0.2000\scriptstyle\pm 0.1732$
L3-FT(QA) $0.4758\scriptstyle\pm 0.0251$ $0.3733\scriptstyle\pm 0.0115$ $0.4180\scriptstyle\pm 0.0082$ $0.2642\scriptstyle\pm 0.0066$ $0.8667\scriptstyle\pm 0.2309$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.1000}$
L3-FT(QA) + KG $\mathbf{0.4988}\scriptstyle\pm\mathbf{0.0700}$ $\mathbf{0.4500}\scriptstyle\pm\mathbf{0.0600}$ $\mathbf{0.4721}\scriptstyle\pm\mathbf{0.0572}$ $\mathbf{0.3102}\scriptstyle\pm\mathbf{0.0481}$ $0.8667\scriptstyle\pm 0.2309$ $0.8333\scriptstyle\pm 0.1155$
G-Retriever $0.3719\scriptstyle\pm 0.0373$ $0.2400\scriptstyle\pm 0.0000$ $0.2912\scriptstyle\pm 0.0119$ $0.1705\scriptstyle\pm 0.0081$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
@@
Table 29: Model performances on DLBC

### v2

@@ E.2 Human evaludation and LLM as judge
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $0053\scriptstyle\pm 0.0092 0.0$$033\scriptstyle\pm 0.0058 0.0$$041\scriptstyle\pm 0.0071 0.0$$21\scriptstyle\pm 0.0036 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0070\scriptstyle\pm 0.0121 0.0$$133\scriptstyle\pm 0.0231 0.0$$092\scriptstyle\pm 0.0159 0.0$$46\scriptstyle\pm 0.0080 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5000\scriptstyle\pm 0.0500$ $0.4367\scriptstyle\pm 0.0702$ $0.4649\scriptstyle\pm 0.0545$ $0.3040\scriptstyle\pm 0.0457$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9667\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics + KG $0.5267\scriptstyle\pm 0.0250$ $0.5067\scriptstyle\pm 0.0115$ $0.5163\scriptstyle\pm 0.0141$ $0.3480\scriptstyle\pm 0.0128$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.1000$
G-Retriever + pre-GAT $0.4846\scriptstyle\pm 0.0537$ $0.4300\scriptstyle\pm 0.0173$ $0.4540\scriptstyle\pm 0.0126$ $0.2937\scriptstyle\pm 0.0106$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
RoG $\mathbf{0.5667}\scriptstyle\pm\mathbf{0.0462}$ $\mathbf{0.5667}\scriptstyle\pm\mathbf{0.0462}$ $\mathbf{0.5667}\scriptstyle\pm\mathbf{0.0462}$ $\mathbf{0.3963}\scriptstyle\pm\mathbf{0.0458}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9667\scriptstyle\pm 0.0577$
SubgraphRAG $0.5400\scriptstyle\pm 0.0000$ $0.5400\scriptstyle\pm 0.0000$ $0.5400\scriptstyle\pm 0.0000$ $0.3699\scriptstyle\pm 0.0000$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
GNN-RAG $0.5294\scriptstyle\pm 0.0000$ $0.2700\scriptstyle\pm 0.0000$ $0.3576\scriptstyle\pm 0.0000$ $0.2177\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
GALAX $0.5433\scriptstyle\pm 0.0252$ $0.5433\scriptstyle\pm 0.0252$ $0.5433\scriptstyle\pm 0.0252$ $0.3733\scriptstyle\pm 0.0238$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
GALAX (Qwen2.5-7B) $0.5501\scriptstyle\pm 0.0576$ $0.5200\scriptstyle\pm 0.0693$ $0.5345\scriptstyle\pm 0.0639$ $0.3664\scriptstyle\pm 0.0581$ $0.8667\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
Table 33: Model performance on NB
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0085\scriptstyle\pm 0.0148 0.0$$67\scriptstyle\pm 0.0289 0.011$$13\scriptstyle\pm 0.0196 0.0$$57\scriptstyle\pm 0.0100 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5597\scriptstyle\pm 0.0479$ $0.5233\scriptstyle\pm 0.0651$ $0.5407\scriptstyle\pm 0.0559$ $0.3718\scriptstyle\pm 0.0523$ $0.8667\scriptstyle\pm 0.2309$ $0.8667\scriptstyle\pm 0.1528$
L3-FT(CRISPR) + Omics + KG $0.5507\scriptstyle\pm 0.0704$ $0.5067\scriptstyle\pm 0.0723$ $0.5275\scriptstyle\pm 0.0703$ $0.3603\scriptstyle\pm 0.0667$ $0.9333\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
G-Retriever + pre-GAT $0.4322\scriptstyle\pm 0.0444$ $0.3933\scriptstyle\pm 0.0404$ $0.4119\scriptstyle\pm 0.0423$ $0.2599\scriptstyle\pm 0.0341$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.0577}$
RoG $\mathbf{0.5997}\scriptstyle\pm\mathbf{0.0551}$ $0.4133\scriptstyle\pm 0.0924$ $0.4817\scriptstyle\pm 0.0400$ $0.3179\scriptstyle\pm 0.0353$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8667\scriptstyle\pm 0.1155$
SubgraphRAG $0.5679\scriptstyle\pm 0.0551$ $0.4667\scriptstyle\pm 0.0924$ $0.5048\scriptstyle\pm 0.0400$ $0.3383\scriptstyle\pm 0.0353$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.1155}$
GNN-RAG $0.5810\scriptstyle\pm 0.0165$ $\mathbf{0.5733}\scriptstyle\pm\mathbf{0.0231}$ $\mathbf{0.5771}\scriptstyle\pm\mathbf{0.0198}$ $\mathbf{0.4058}\scriptstyle\pm\mathbf{0.0197}$ $0.7333\scriptstyle\pm 0.2309$ $0.7333\scriptstyle\pm 0.0577$
GALAX $0.5533\scriptstyle\pm 0.0115$ $0.5533\scriptstyle\pm 0.0115$ $0.5533\scriptstyle\pm 0.0115$ $0.3825\scriptstyle\pm 0.0110$ $0.8667\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
GALAX (Qwen2.5-7B) $0.5733\scriptstyle\pm 0.0462$ $\mathbf{0.5733}\scriptstyle\pm\mathbf{0.0462}$ $0.5733\scriptstyle\pm 0.0462$ $0.4028\scriptstyle\pm 0.0446$ $0.8000\scriptstyle\pm 0.0000$ $0.8667\scriptstyle\pm 0.0577$
Table 34: Model performance on MESO
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0190\scriptstyle\pm 0.0201 0.0$$167\scriptstyle\pm 0.0208 0.0$$133\scriptstyle\pm 0.0121 0.0$$67\scriptstyle\pm 0.0061 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0192\scriptstyle\pm 0.0228 0.0$$100\scriptstyle\pm 0.0000 0.0$$104\scriptstyle\pm 0.0052 0.0$$53\scriptstyle\pm 0.0026 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5182\scriptstyle\pm 0.0806$ $0.4833\scriptstyle\pm 0.0874$ $0.4995\scriptstyle\pm 0.0808$ $0.3354\scriptstyle\pm 0.0720$ $0.9333\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.1528$
L3-FT(CRISPR) + Omics + KG $0.5417\scriptstyle\pm 0.0419$ $0.5233\scriptstyle\pm 0.0416$ $0.5322\scriptstyle\pm 0.0402$ $0.3633\scriptstyle\pm 0.0374$ $0.8667\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
G-Retriever + pre-GAT $0.4721\scriptstyle\pm 0.0072$ $0.3733\scriptstyle\pm 0.0462$ $0.4161\scriptstyle\pm 0.0324$ $0.2630\scriptstyle\pm 0.0255$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8000\scriptstyle\pm 0.0000$
RoG $0.4476\scriptstyle\pm 0.0755$ $0.3900\scriptstyle\pm 0.1375$ $0.4137\scriptstyle\pm 0.1135$ $0.2651\scriptstyle\pm 0.0893$ $0.8667\scriptstyle\pm 0.2309$ $0.8000\scriptstyle\pm 0.0000$
SubgraphRAG $0.4421\scriptstyle\pm 0.0000$ $0.4200\scriptstyle\pm 0.0000$ $0.4308\scriptstyle\pm 0.0000$ $0.2745\scriptstyle\pm 0.0000$ $0.6000\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
GNN-RAG $0.3750\scriptstyle\pm 0.0000$ $0.2400\scriptstyle\pm 0.0000$ $0.2927\scriptstyle\pm 0.0000$ $0.1714\scriptstyle\pm 0.0000$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8000\scriptstyle\pm 0.0000$
@@
GALAX (Qwen2.5-7B) $0.5621\scriptstyle\pm 0.0310$ $0.5533\scriptstyle\pm 0.0462$ $0.5576\scriptstyle\pm 0.0388$ $0.3872\scriptstyle\pm 0.0367$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.0577}$
Table 35: Model performance on LIHC
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0031\scriptstyle\pm 0.0027 0.0$$067\scriptstyle\pm 0.0058 0.0$$043\scriptstyle\pm 0.0037 0.0$$21\scriptstyle\pm 0.0019 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.4239\scriptstyle\pm 0.0343$ $0.3967\scriptstyle\pm 0.0513$ $0.4096\scriptstyle\pm 0.0435$ $0.2582\scriptstyle\pm 0.0341$ $0.7333\scriptstyle\pm 0.1155$ $0.7000\scriptstyle\pm 0.1732$
L3-FT(CRISPR) + Omics + KG $0.4552\scriptstyle\pm 0.0568$ $0.4467\scriptstyle\pm 0.0635$ $0.4509\scriptstyle\pm 0.0600$ $0.2924\scriptstyle\pm 0.0512$ $0.8000\scriptstyle\pm 0.0000$ $0.7667\scriptstyle\pm 0.0577$
G-Retriever + pre-GAT $0.3598\scriptstyle\pm 0.0387$ $0.3000\scriptstyle\pm 0.0520$ $0.3270\scriptstyle\pm 0.0467$ $0.1961\scriptstyle\pm 0.0340$ $0.8000\scriptstyle\pm 0.0000$ $0.7667\scriptstyle\pm 0.1155$
RoG $0.4205\scriptstyle\pm 0.1173$ $0.3733\scriptstyle\pm 0.1856$ $0.3903\scriptstyle\pm 0.1596$ $0.2507\scriptstyle\pm 0.1247$ $0.9333\scriptstyle\pm 0.1155$ $0.8000\scriptstyle\pm 0.1732$
SubgraphRAG $0.4738\scriptstyle\pm 0.1320$ $0.4267\scriptstyle\pm 0.2136$ $0.4436\scriptstyle\pm 0.1843$ $0.2964\scriptstyle\pm 0.1437$ $0.9333\scriptstyle\pm 0.1155$ $0.8000\scriptstyle\pm 0.1732$
GNN-RAG $0.4233\scriptstyle\pm 0.0289$ $0.4233\scriptstyle\pm 0.0289$ $0.4233\scriptstyle\pm 0.0289$ $0.2688\scriptstyle\pm 0.0230$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8333\scriptstyle\pm 0.0577$
@@
GALAX (Qwen2.5-7B) $0.4362\scriptstyle\pm 0.0033$ $0.4333\scriptstyle\pm 0.0058$ $0.4348\scriptstyle\pm 0.0045$ $0.2778\scriptstyle\pm 0.0037$ $0.8667\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
Table 36: Model performance on LAML
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $1576\scriptstyle\pm 0.1412 0.0$$200\scriptstyle\pm 0.0173 0.0$$354\scriptstyle\pm 0.0307 0.0$$0182\scriptstyle\pm 0.0158$ $0.1333\scriptstyle\pm 0.2309$ $0.2000\scriptstyle\pm 0.1732$
L3-FT(Med) + Omics $1136\scriptstyle\pm 0.0630 0.0$$400\scriptstyle\pm 0.0173 0.0$$495\scriptstyle\pm 0.0008 0.0$$0254\scriptstyle\pm 0.0004$ $0.2000\scriptstyle\pm 0.3464$ $0.1333\scriptstyle\pm 0.1528$
L3-FT(Med) + Omics + KG $1085\scriptstyle\pm 0.0867 0.0$$267\scriptstyle\pm 0.0058 0.0$$377\scriptstyle\pm 0.0221 0.0$$0193\scriptstyle\pm 0.0114$ $0.2000\scriptstyle\pm 0.3464$ $0.2000\scriptstyle\pm 0.1732$
L3-FT(CRISPR) + Omics $0.4758\scriptstyle\pm 0.0251$ $0.3733\scriptstyle\pm 0.0115$ $0.4180\scriptstyle\pm 0.0082$ $0.2642\scriptstyle\pm 0.0066$ $0.8667\scriptstyle\pm 0.2309$ $0.9000\scriptstyle\pm 0.1000$
L3-FT(CRISPR) + Omics + KG $\mathbf{0.4988}\scriptstyle\pm\mathbf{0.0700}$ $0.4500\scriptstyle\pm 0.0600$ $\mathbf{0.4721}\scriptstyle\pm\mathbf{0.0572}$ $\mathbf{0.3102}\scriptstyle\pm\mathbf{0.0481}$ $0.8667\scriptstyle\pm 0.2309$ $0.8333\scriptstyle\pm 0.1155$
G-Retriever + pre-GAT $0.3719\scriptstyle\pm 0.0373$ $0.2400\scriptstyle\pm 0.0000$ $0.2912\scriptstyle\pm 0.0119$ $0.1705\scriptstyle\pm 0.0081$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
RoG $0.4228\scriptstyle\pm 0.0149$ $0.3467\scriptstyle\pm 0.1097$ $0.3725\scriptstyle\pm 0.0685$ $0.2303\scriptstyle\pm 0.0506$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8667\scriptstyle\pm 0.1155$
SubgraphRAG $0.4141\scriptstyle\pm 0.0000$ $0.4100\scriptstyle\pm 0.0000$ $0.4121\scriptstyle\pm 0.0000$ $0.2595\scriptstyle\pm 0.0000$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8000\scriptstyle\pm 0.0000$
GNN-RAG $0.3972\scriptstyle\pm 0.0934$ $0.3200\scriptstyle\pm 0.1559$ $0.3511\scriptstyle\pm 0.1311$ $0.2183\scriptstyle\pm 0.1015$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8333\scriptstyle\pm 0.0577$
@@
GALAX (Qwen2.5-7B) $0.4596\scriptstyle\pm 0.0437$ $\mathbf{0.4567}\scriptstyle\pm\mathbf{0.0462}$ $0.4581\scriptstyle\pm 0.0449$ $0.2979\scriptstyle\pm 0.0385$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
Table 37: Model performance on DLBC